Parth (parthkohli):

Pop quiz for y'all minds out there!\[\large \sum_{i = 2}^{25}P(i,2)\]

Parth (parthkohli):

Ssh. ;)

hartnn (hartnn):

P=permutation

OpenStudy (anonymous):

I'm not familiar with this notation, what is \(P(i, 2)\) representing?

OpenStudy (anonymous):

oh.

hartnn (hartnn):

i won't just clearing what's P

Parth (parthkohli):

OK, thank you!

OpenStudy (anonymous):

\[=\sum{\frac{i!}{(i-2)!}}=\sum{i(i-1)}=\sum{(i^2-i)}\] for i=1 to 25. \[\sum{i^2}=\frac{n(n+1)(2n+1)}{6}\] For n=25 this is 5525. \[\sum{i}=\frac{n(n+1)}{2}\] For n=25 this is 325. So \(\sum{(i^2-i)}=5525-325=5200\) Did I go wrong somewhere?

hartnn (hartnn):

wow!

Parth (parthkohli):

Very good, very good!

hartnn (hartnn):

in 1st shot!

OpenStudy (anonymous):

Yay, thanks :)